OpenStudy (anonymous):

Amy paid $260 for an airplane ticket plus a sales tax of 23% What is the total cost that Amy Paid?

Directrix (directrix):

260 + .23*(260) = or 1.23 * (260) =

OpenStudy (asimraza):

\[{260+(23.260)/100}=319\] answer is 319